# Vision Transformers in Precision Agriculture: A Comprehensive Survey

### v3

@@ Abstract
Detecting plant diseases is a crucial aspect of modern agriculture, as it plays a key role in maintaining crop health and increasing overall yield. Traditional approaches, though still valuable, often rely on manual inspection or conventional machine learning techniques, both of which face limitations in scalability and accuracy. Recently, Vision Transformers (ViTs) have emerged as a promising alternative, offering advantages such as improved handling of long-range dependencies and better scalability for visual tasks. This review explores the application of ViTs in precision agriculture, covering a range of tasks. We begin by introducing the foundational architecture of ViTs and discussing their transition from Natural Language Processing (NLP) to Computer Vision. The discussion includes the concept of inductive bias in traditional models like Convolutional Neural Networks (CNNs), and how ViTs mitigate these biases. We provide a comprehensive review of recent literature, focusing on key methodologies, datasets, and performance metrics. This study also includes a comparative analysis of CNNs and ViTs, along with a review of hybrid models and performance enhancements. Technical challenges such as data requirements, computational demands, and model interpretability are addressed, along with potential solutions. Finally, we outline future research directions and technological advancements that could further support the integration of ViTs in real-world agricultural settings. Our goal with this study is to offer practitioners and researchers a deeper understanding of how ViTs are poised to transform smart and precision agriculture.
@@ 1 Introduction
Plant diseases pose a significant challenge to global agriculture, affecting crop yields, food security, and the economic well-being of farmers. Early and accurate detection is essential to effectively manage these diseases and minimize their impact [11]. Traditionally, the diagnosis of plant diseases has depended on labor-intensive manual inspections by experts or on conventional Machine Learning (ML) techniques, both of which have inherent limitations.Manual inspection is not only time-consuming but also subject to human error and inconsistency, making it impractical for large-scale agricultural operations [57, 26]. Conventional ML methods are largely dependent on manual feature extraction, which limits their capacity to generalize across diverse environmental conditions, such as variations in lighting, background, and plant species. Moreover, these techniques are unable to effectively capture complex spatial patterns directly from raw image data. [24, 59]. To address these limitations, researchers are increasingly turning to automated solutions that harness recent advances in ML and computer vision. CNN, a widely used traditional Deep Learning (DL) method, have shown some success in automating the disease detection process [90, 52]. This has created a growing demand for more robust, scalable, and intelligent systems for plant disease detection. In response, several studies have explored the potential of DL to meet these challenges. Some existing surveys provide general overviews of DL approaches, highlighting their advantages over traditional methods and outlining the progress made in applying these techniques to plant disease detection tasks. For example, one study [85] reviews how DL and ML techniques have been employed for plant disease detection from image data. The study describes the transition from handcrafted features with the traditional DL methods to deep models like CNNs, DBNs, and segmentation models. It also discusses the implementation of object detection algorithms like YOLO and Faster R-CNN, and segmentation architectures like U-Net and Mask R-CNN, for pixel-level accuracy-based detection quality improvement. Another work [42] discusses the application of ML and DL techniques in the early detection and classification of plant diseases, highlighting their importance in improving agricultural yield and reducing crop loss. It presents an extensive list of ML techniques such as Naïve Bayes, k-Nearest Neighbors, Decision Trees, Support Vector Machines, Random Forests, and Multi-layer Perceptrons, and explores their applications in plant disease classification using environmental and image inputs. The study also discusses preprocessing techniques and image segmentation methods, demonstrating how ML has enabled the automation of disease diagnosis and soil testing from models trained with crop data. The contrast between classical ML and DL methods shows that DL outperforms ML in terms of accuracy and scalability. This study emphasizes the importance of addressing current deficiencies and proposes DL as a promising direction for creating automatic, systematic, and scalable plant disease diagnostic systems that support both scientists and farmers in agricultural decision-making. Another overview of studies [2] explores the use of DL techniques for plant disease detection and management in agriculture. It addresses key challenges including dataset quality, imaging sensors, model generalizability, estimation of severity, and performance comparison with human capabilities. Commonly used datasets are categorized, along with their limitations in practical scenarios. The study outlines various DL techniques, with a particular focus on CNN-based models, employed for image classification, object detection, and semantic segmentation. It highlights limitations in the ability of existing models to generalize to field conditions, primarily due to overfitting and imbalanced data. Additionally, significant research areas are highlighted, such as universal severity estimation, multi-disease detection, and model generalization across diverse conditions. A separate study [53] explores the application of DL in plant disease recognition and classification, emphasizing its critical role in agricultural productivity and food security. It illustrates how DL improves upon traditional image processing techniques through automated feature extraction and enhanced objectivity and efficiency. The paper discusses foundational concepts, model performance criteria, and the evolution of DL in plant disease recognition. It also addresses the need for data augmentation to manage limited datasets and the use of visualization methods to improve model interpretability. The study concludes with future challenges, such as the need for more varied datasets, enhanced model robustness, and the integration of hyperspectral imaging for early detection.
@@
Recent advancements in deep learning have demonstrated the potential of CNNs and their variants in accurately detecting and classifying plant diseases. However, despite their success, CNN-based models often struggle to capture global contextual information due to their inherent inductive biases and locality constraints. The introduction of the Transformer architecture has transformed the field of NLP, enabling significant advancements in machine translation, text summarization, and language modeling, and setting new performance benchmarks across a wide range of language understanding tasks [69]. By leveraging parallel processing and self-attention mechanisms, Transformers excel at capturing long-range dependencies and contextual relationships. This breakthrough has inspired researchers to explore the application of Transformers in other domains, including Computer Vision. It was soon recognized that the same properties that made Transformers effective in NLP, namely their ability to model long-range dependencies and contextually relevant information, could also be advantageous for image analysis tasks [21]. Unlike CNNs, which rely on inductive biases such as local connectivity and spatial hierarchies, Transformers offer a more flexible architecture with fewer assumptions about input structure. This flexibility enables them to learn rich, complex patterns directly from visual data, making them a promising alternative for various computer vision applications, including automated plant disease detection.
@@
The use of ViTs in agriculture is a novel and rapidly emerging research area. To the best of our knowledge, there are a few works whose focus is on ViTs; however, most of them mention it only briefly and do not provide a comprehensive analysis of their applications in precision agriculture. For example, a review [63] categorizes DL techniques applied to plant disease detection, organizing the literature into three tasks: classification, detection, and segmentation of plant diseases from leaf images. The review evaluates a broad range of DL models, with a strong focus on CNNs such as ResNet, VGG, and EfficientNet, alongside growing interest in ViTs. Another study [103] examines the use of DL and Computer Vision in plant disease detection, systematically analyzing the transition from traditional image processing to modern DL-based approaches. It assesses the suitability and performance of image processing techniques like thresholding, edge detection, and region-based segmentation in real-world agriculture. The study categorizes various DL architectures, such as CNNs, Generative Adversarial Network (GAN), ViTs, and Vision-Language Models, and highlights their scalability and potential for practical implementation. It covers advanced models like YOLO and Faster R-CNN for real-time object detection, and discusses the emerging relevance of ViTs for capturing global context in image analysis. However, while ViTs are introduced, the review lacks a deep evaluation of specific architectures or performance benchmarks in agricultural contexts.
@@
In this work, we examine the use of ViT in the agriculture domain with the objective of identifying how this architecture has been used and adapted for agricultural applications such as crop classification, disease detection, yield prediction, and precision agriculture. With ViT’s success in Computer Vision across numerous domains, its application to transforming agricultural image analysis holds much promise. By analyzing the existing literature, we aspire to extract current trends, evaluate model performance, and identify challenges specific to agricultural environments and datasets. This initial exploration will lay the groundwork for future studies and build a sound body of knowledge within this emerging interface of DL and agriculture.
@@
  • An overview of the Transformer and Vision Transformer architectures and their functionality is provided.
@@ 2.1 Transformer
The Transformer model, proposed by [104], brought a groundbreaking shift to NLP. Its core innovation is the self-attention mechanism, which enables the model to assess the significance of various words within a sentence during sequence encoding and generation. This mechanism addresses the limitations of earlier architectures, such as Recurrent Neural Networks (RNNs) and CNNs, which often struggled with long-range dependencies and parallelization challenges. By processing all tokens in a sequence simultaneously, self-attention enables the Transformer to model long-range dependencies more effectively, making the architecture highly efficient and scalable for various sequence tasks. In the following subsection, we examine its main components.
@@ 2.1.1 Encoder-Decoder structure
The Transformer consists of an encoder and a decoder, both of which are composed of multiple layers of Self-Attention and Feed-Forward Neural Networks. The encoder processes the input sequence, while the decoder generates the output sequence. This architecture is typically used in tasks like machine translation, although variations of the Transformer, such as GPT [10], use only the Decoder for language modeling. Figure 1 illustrates the Encoder-Decoder structure.
@@
Each encoder layer consists of two main components:
@@
These components are followed by residual connections and layer normalization. Each decoder layer includes:
@@
Similar to the encoder, these components are followed by residual connections and layer normalization.
@@ 2.1.3 Multi-Head Attention
Figure 3 shows the multi-head attention mechanism.
@@ 2.2 Vision Transformer
Inspired by the success of Transformers in NLP, [21] adapted the Transformer architecture for computer vision tasks, leading to the creation of ViT. Before this breakthrough, visual attention mechanisms were typically combined with convolutional networks or used to enhance specific components, while retaining the overall CNN structure. Early attention-based architectures had shown promise in computer vision [40], but their implementation on hardware accelerators often required complex and specialized engineering, as these systems were primarily optimized for convolutional operations. The key insight of ViT was to show that convolutional networks are not essential: a simple Transformer applied directly to sequences of image patches can achieve strong results in image classification. In ViT, images are divided into fixed-size patches, each treated as a token, similar to words in NLP. These patches are linearly embedded, combined with positional embeddings, and processed through multiple Self-Attention layers. This design enables ViTs to effectively capture long-range dependencies and contextual relationships within images, leading to performance that often surpasses traditional CNNs on various classification benchmarks. By leveraging the Self-Attention mechanism of Transformers, ViTs have revolutionized the field of computer vision. The following section delves into the architecture of ViTs, explaining their key components and operations in detail. Figure 4 shows the structure of ViT. The architecture of ViT consists of the following key components: Patch Embedding, Positional Embedding, Transformer Encoder, and Classification Head.
@@ 2.2.1 Patch Embedding
where $z^{P}_{0}\in\mathbb{R}^{D}$ is the embedding of the $P^{th}$ patch at Transformer layer-00 (before any attention layers), $x^{P}$ is the $P^{th}$ image patch before projection and $E\in\mathbb{R}^{(P^{2}.C)\times D}$ is a learnable weight matrix that projects each flattened image patch into a $D$-dimensional embedding space.
@@ 2.2.2 Positional Embedding
where $z_{0}$ is the full embedded input sequence after adding positional information at Transformer layer-00 (before any attention layers), $x_{i}\in\mathbb{R}^{P^{2}.C}$ is the $i^{th}$ flattened image patch, $E\in\mathbb{R}^{(P^{2}.C)\times D}$ is the patch embedding projection matrix and $P\in\mathbb{R}^{N\times D}$ is the positional embeddings. The positional embeddings added to the patch embeddings can be either learned or fixed.
@@ 2.2.3 Transformer Encoder
where $z_{l-1}$ is the output from the previous encoder layer, $MHSA(z_{l-1})$ is the result of applying multi-head self-attention to that output, $z^{{}^{\prime}}_{l}$ is the input to the feed-forward network (FFN) sub-layer.$z_{l}$ denotes the output of $l^{th}$layer.
@@ 3.1 Definition of Inductive Biases
In supervised learning, when the data is frequently incomplete or noisy, inductive biases are essential for the model’s effectiveness. In Computer Vision, for example in CNNs, some of the inherent inductive biases include locality and translation equivariance [20, 72, 25]. Locality assumes that important features, such as edges and textures, are restricted to small spatial regions within an image, while translation equivariance guarantees that these features are identified regardless of their position in the image. These biases enable CNNs to perform exceptionally well in tasks like image classification and object detection, where spatial hierarchies play a key role [47, 73]. Similarly, in NLP, RNNs and Transformers utilize sequential and contextual relationships as biases to process and understand text. The type of inductive biases has a direct impact on a model’s performance on particular tasks.
@@ 3.3 Strategies to Mitigate the Lack of Inductive Biases in ViTs
  • Convolutional Patch Embedding: Replacing simple patch embedding with convolutional layers to preserve local structures in the image, or employing recursive token aggregation, which functions similarly to convolutions in capturing local spatial relationships, is a method utilized in models such as Tokens-to-Token ViT (T2T-ViT) [113].
  • Hybrid Models: As mentioned earlier, CNNs are effective at capturing local dependencies, while ViTs excel at modeling long-range global dependencies. One way to mitigate the lack of inductive biases in ViTs is to use a CNN or convolutional backbone to extract low-level features, then feed this output into a ViT. This approach combines the spatial inductive biases of CNNs with the global modeling capabilities of ViTs. Introducing a few convolutional inductive biases in the early stages of ViT processing can strike a balance between inductive bias and the learning capacity of Transformer blocks, while also influencing the optimization behavior of ViTs [108].
  • Knowledge Distillation: Distilling knowledge that comes from CNN into ViTs during training is another possible solution. As an example, DEiT uses this technique [101]. It uses a CNN as the teacher and a ViT as the student to improve its performance by learning from the teacher during the training process

### v4

@@ Abstract
Detecting plant diseases is a crucial aspect of modern agriculture, as it plays a key role in maintaining crop health and increasing overall yield. Traditional approaches, though still valuable, often rely on manual inspection or conventional machine learning techniques, both of which face limitations in scalability and accuracy. Recently, Vision Transformers (ViTs) have emerged as a promising alternative, offering advantages such as improved handling of long-range dependencies and better scalability for visual tasks. This review explores the application of ViTs in precision agriculture, covering a range of tasks. We begin by introducing the foundational architecture of ViTs and discussing their transition from Natural Language Processing (NLP) to Computer Vision (CV). The discussion includes the concept of inductive bias in traditional models like Convolutional Neural Networks (CNNs), and how ViTs mitigate these biases. We provide a comprehensive review of recent literature, focusing on key methodologies, datasets, and performance metrics. This study also includes a comparative analysis of CNNs and ViTs, along with a review of hybrid models and performance enhancements. Technical challenges such as data requirements, computational demands, and model interpretability are addressed, along with potential solutions. Finally, we outline future research directions and technological advancements that could further support the integration of ViTs in real-world agricultural settings. Our goal with this study is to offer practitioners and researchers a deeper understanding of how ViTs are poised to transform smart and precision agriculture.
@@ 1 Introduction
Plant diseases pose a significant challenge to global agriculture, affecting crop yields, food security, and the economic well-being of farmers. Early and accurate detection is essential to effectively manage these diseases and minimize their impact [11]. Traditionally, the diagnosis of plant diseases has depended on labor-intensive manual inspections by experts or on conventional Machine Learning (ML) techniques, both of which have inherent limitations. Manual inspection is not only time-consuming but also subject to human error and inconsistency, making it impractical for large-scale agricultural operations [58, 26]. Conventional ML methods are largely dependent on manual feature extraction, which limits their capacity to generalize across diverse environmental conditions, such as variations in lighting, background, and plant species. Moreover, these techniques are unable to effectively capture complex spatial patterns directly from raw image data [24, 60]. To address these limitations, researchers are increasingly turning to automated solutions that harness recent advances in ML and CV.
@@
CNN, a widely used traditional Deep Learning (DL) method, have shown some success in automating the disease detection process [91, 53]. This has created a growing demand for more robust, scalable, and intelligent systems for plant disease detection. In response, several studies have explored the potential of DL to meet these challenges. Some existing surveys provide general overviews of DL approaches, highlighting their advantages over traditional methods and outlining the progress made in applying these techniques to plant disease detection tasks. For example, one study [86] reviews how DL and ML techniques have been employed for plant disease detection from image data. The study describes the transition from handcrafted features with the traditional DL methods to deep models like CNNs and segmentation models. It also discusses the implementation of object detection algorithms like YOLO and Faster R-CNN, and segmentation architectures like U-Net and Mask R-CNN, for pixel-level accuracy-based detection quality improvement. Another work [43] discusses the application of ML and DL techniques in the early detection and classification of plant diseases, highlighting their importance in improving agricultural yield and reducing crop loss. It presents an extensive list of ML techniques such as Naïve Bayes, k-Nearest Neighbors, Decision Trees, Support Vector Machines, Random Forests, and Multi-layer Perceptrons, and explores their applications in plant disease classification using environmental and image inputs. The study also discusses preprocessing techniques and image segmentation methods, demonstrating how ML has enabled the automation of disease diagnosis and soil testing from models trained with crop data. The contrast between classical ML and DL methods shows that DL outperforms ML in terms of accuracy and scalability. This study emphasizes the importance of addressing current deficiencies and proposes DL as a promising direction for creating automatic, systematic, and scalable plant disease diagnostic systems that support both scientists and farmers in agricultural decision-making. Another overview of studies [3] explores the use of DL techniques for plant disease detection and management in agriculture. It addresses key challenges including dataset quality, imaging sensors, model generalizability, estimation of severity, and performance comparison with human capabilities. Commonly used datasets are categorized, along with their limitations in practical scenarios. The study outlines various DL techniques, with a particular focus on CNN-based models, employed for image classification, object detection, and semantic segmentation. It highlights limitations in the ability of existing models to generalize to field conditions, primarily due to overfitting and imbalanced data. Additionally, significant research areas are highlighted, such as universal severity estimation, multi-disease detection, and model generalization across diverse conditions. A separate study [54] explores the application of DL in plant disease recognition and classification, emphasizing its critical role in agricultural productivity and food security. It illustrates how DL improves upon traditional image processing techniques through automated feature extraction and enhanced objectivity and efficiency. The paper discusses foundational concepts, model performance criteria, and the evolution of DL in plant disease recognition. It also addresses the need for data augmentation to manage limited datasets and the use of visualization methods to improve model interpretability. The study concludes with future challenges, such as the need for more varied datasets, enhanced model robustness, and the integration of hyperspectral imaging for early detection.
@@
Recent advancements in DL have demonstrated the potential of CNNs and their variants in accurately detecting and classifying plant diseases. However, despite their success, CNN-based models often struggle to capture global contextual information due to their inherent inductive biases and locality constraints. The introduction of the Transformer architecture has transformed the field of NLP, enabling significant advancements in machine translation, text summarization, and language modeling, and setting new performance benchmarks across a wide range of language understanding tasks [70]. By leveraging parallel processing and self-attention mechanisms, Transformers excel at capturing long-range dependencies and contextual relationships. This breakthrough has inspired researchers to explore the application of Transformers in other domains, including CV. It was soon recognized that the same properties that made Transformers effective in NLP, namely their ability to model long-range dependencies and contextually relevant information, could also be advantageous for image analysis tasks [21]. Unlike CNNs, which rely on inductive biases such as local connectivity and spatial hierarchies, Transformers offer a more flexible architecture with fewer assumptions about input structure. This flexibility enables them to learn rich, complex patterns directly from visual data, making them a promising alternative for various CV applications, including automated plant disease detection.
@@
The use of ViTs in agriculture is a novel and rapidly emerging research area. To the best of our knowledge, there are a few works whose focus is on ViTs; however, most of them mention it only briefly and do not provide a comprehensive analysis of their applications in precision agriculture. For example, a review [64] categorizes DL techniques applied to plant disease detection, organizing the literature into three tasks: classification, detection, and segmentation of plant diseases from leaf images. The review evaluates a broad range of DL models, with a strong focus on CNNs such as ResNet [33], VGG [88], and EfficientNet [94], alongside growing interest in ViTs. Another study [104] examines the use of DL and CV in plant disease detection, systematically analyzing the transition from traditional image processing to modern DL-based approaches. It assesses the suitability and performance of image processing techniques like thresholding, edge detection, and region-based segmentation in real-world agriculture. The study categorizes various DL architectures, such as CNNs, Generative Adversarial Network (GAN), ViTs, and Vision-Language Models, and highlights their scalability and potential for practical implementation. It covers advanced models like YOLO and Faster R-CNN for real-time object detection, and discusses the emerging relevance of ViTs for capturing global context in image analysis. However, while ViTs are introduced, the review lacks a deep evaluation of specific architectures or performance benchmarks in agricultural contexts.
@@
In this work, we examine the use of ViT in the agriculture domain with the objective of identifying how this architecture has been used and adapted for agricultural applications such as crop classification, disease detection, yield prediction, and precision agriculture. With ViT’s success in CV across numerous domains, its application to transforming agricultural image analysis holds much promise. By analyzing the existing literature, we aspire to extract current trends, evaluate model performance, and identify challenges specific to agricultural environments and datasets. This initial exploration will lay the groundwork for future studies and build a sound body of knowledge within this emerging interface of DL and agriculture.
@@
  • An overview of the Transformer and ViT architectures and their functionality is provided.
@@ 2.1 Transformer
The Transformer model [105] brought a groundbreaking shift to NLP by introducing the self-attention mechanism, which allows the model to evaluate the importance of different words in a sentence during sequence encoding and generation. This innovation overcomes the limitations of earlier architectures, such as Recurrent Neural Networks (RNNs) and CNNs, which often struggled with long-range dependencies and parallelization. By processing all tokens in a sequence simultaneously, self-attention captures long-range relationships more effectively, making the Transformer architecture both efficient and highly scalable for a wide range of sequence modeling tasks. In the following subsection, we examine its main components.
@@ 2.1.1 Encoder-Decoder structure
The Transformer architecture comprises an encoder and a decoder, each consisting of multiple layers of self-attention and Feed-Forward Neural Networks. The encoder processes the input sequence into a contextual representation, while the decoder generates the output sequence based on both the encoder’s output and previously generated tokens. This encoder–decoder structure is commonly used in tasks such as machine translation, whereas certain variants, such as GPT [10], employ only the decoder for autoregressive language modeling. Figure 1 illustrates the overall architecture of the Transformer’s encoder–decoder structure.
@@
Each encoder layer consists of two main sublayers:
@@
Each sublayer is followed by a residual connection and layer normalization.
Each decoder layer consists of three sublayers:
@@
Similar to the encoder, each sublayer is followed by a residual connection and layer normalization.
@@ 2.1.3 Multi-Head Attention
Figure 3 provides a schematic representation of the multi-head attention mechanism.
@@ 2.2 Vision Transformer
Inspired by the success of Transformers in NLP, A. Dosovitskiy et al. [21] adapted the Transformer architecture for CV tasks, leading to the creation of ViT. Before this breakthrough, visual attention mechanisms were typically combined with convolutional networks or used to enhance specific components, while retaining the overall CNN structure. Early attention-based architectures had shown promise in CV [41], but their implementation on hardware accelerators often required complex and specialized engineering, as these systems were primarily optimized for convolutional operations. The key insight of ViT was to show that convolutional networks are not essential: a simple Transformer applied directly to sequences of image patches can achieve strong results in image classification. In ViT, images are divided into fixed-size patches, each treated as a token, similar to words in NLP. These patches are linearly embedded, combined with positional embeddings, and processed through multiple Self-Attention layers. This design enables ViTs to effectively capture long-range dependencies and contextual relationships within images, leading to performance that often surpasses traditional CNNs on various classification benchmarks. By leveraging the Self-Attention mechanism of Transformers, ViTs have revolutionized the field of CV. The following section delves into the architecture of ViTs, explaining their key components and operations in detail. Figure 4 shows the structure of ViT. The architecture of ViT consists of the following key components: Patch Embedding, Positional Embedding, Transformer Encoder, and Classification Head.
@@ 2.2.1 Patch Embedding
where $z^{P}_{0}\in\mathbb{R}^{D}$ is the embedding of the $P^{th}$ patch at Transformer layer-0 (before any attention layers), $x^{P}$ is the $P^{th}$ image patch before projection and $E\in\mathbb{R}^{(P^{2}.C)\times D}$ is a learnable weight matrix that projects each flattened image patch into a $D$-dimensional embedding space.
@@ 2.2.2 Positional Embedding
where $z_{0}$ is the full embedded input sequence after adding positional information at Transformer layer-0 (before any attention layers), $x_{i}\in\mathbb{R}^{P^{2}.C}$ is the $i^{th}$ flattened image patch, $E\in\mathbb{R}^{(P^{2}.C)\times D}$ is the patch embedding projection matrix and $P\in\mathbb{R}^{N\times D}$ is the positional embeddings. The positional embeddings added to the patch embeddings can be either learned or fixed.
@@ 2.2.3 Transformer Encoder
[…]
@@ 3.1 Definition of Inductive Biases
In supervised learning, when the data is frequently incomplete or noisy, inductive biases are essential for the model’s effectiveness. InCV, for example in CNNs, some of the inherent inductive biases include locality and translation equivariance [20, 73, 25]. Locality assumes that important features, such as edges and textures, are restricted to small spatial regions within an image, while translation equivariance guarantees that these features are identified regardless of their position in the image. These biases enable CNNs to perform exceptionally well in tasks like image classification and object detection, where spatial hierarchies play a key role [49, 74]. Similarly, in NLP, RNNs and Transformers utilize sequential and contextual relationships as biases to process and understand text. The type of inductive biases has a direct impact on a model’s performance on particular tasks.
@@ 3.3 Strategies to Mitigate the Lack of Inductive Biases in ViTs
  • Convolutional Patch Embedding: It involves replacing simple patch embedding with convolutional layers to preserve local structures in the image, or employing recursive token aggregation, which functions similarly to convolutions in capturing local spatial relationships. This strategy is utilized in models such as Tokens-to-Token ViT (T2T-ViT) [114].
  • Hybrid Models: One way to mitigate the lack of inductive biases in ViTs is to use a CNN or convolutional backbone to extract low-level features and then feed this output into a ViT. This approach combines the spatial inductive biases of CNNs with the global modeling capabilities of ViTs. Introducing a few convolutional inductive biases in the early stages of ViT processing can strike a balance between inductive bias and the learning capacity of Transformer blocks, while also influencing the optimization behavior of ViTs [109].
  • Knowledge Distillation: Distilling knowledge from CNNs into ViTs during training is another possible solution, as exemplified by DEiT [101], which uses a CNN as the teacher and a ViT as the student to improve performance through teacher–student learning during the training process.